# Simulating stochastic population dynamics: The Linear Noise Approximation can capture non-linear phenomena

### v3

@@ 2.5.5 Generalised pcLNA
2.5.6 Defining the phase correction map
Recall that we wish to define the phase correction map G{G} to take any stochastic state $\bm{X}$} and map it to a point on a deterministic evolution of the system described in (6). The results in Section 2.5.4 suggest phase correction maps that are specific to each class of qualitatively identical (that is, l.t.e.) dynamics, and apply phase correction on the non-linear coordinates of the system using a suitable map for the type of dynamics presented on exemplar systems in the class.
This translates to the following general steps for defining a phase correction map $G$ for any stochastic state $\bm{X}$}:
@@ 2.5.6 Defining the phase correction map
    Suppose that a RRE system as in (6) has a non-hyperbolic equilibrium, $({\bm{x}}_{\alpha_{p}},\alpha_{p})$. That is, ${\bm{F}}({\bm{x}}_{\alpha_{p}},\alpha_{p})={\bm{0}}$, and the Jacobian matrix $\bm{J}({\bm{x}}_{\alpha_{p}},\alpha_{p})$ has $n_{0}$ ($n_{0}>0$) eigenvalues with zero real part.
@@
    Identify the class of systems that the RRE system belongs to by checking whether it satisfies certain conditions such as those described in Sections 3 and 4.
@@
    For $\alpha$ in the vicinity of $\alpha_{p}$, solve the RRE system to derive trajectories $\gamma^{(j)}=\{{\bm{x}}^{(j)}(t):t\geq 0\}$, $j=1,2,\dots,J$.
@@
    Define a projection map $\text{proj}:\mathbb{R}^{n}\to\mathbb{R}^{n_{0}}$, ${\bm{u}}=\text{proj}({\bm{x}})={\bm{R}}_{1}^{T}({\bm{x}}-{\bm{x}}_{\alpha})$, where ${\bm{R}}_{1}$ a $n\times n_{0}$ matrix with columns an eigenbasis corresponding to the $n_{0}$ eigenvalues of $\bm{J}({\bm{x}}_{\alpha},\alpha)$ that for $\alpha=\alpha_{p}$ have zero real part.
@@
    Compute $\{{\bm{u}}^{(j)}(t):t\geq 0\}$, where ${\bm{u}}^{(j)}(t)=\text{proj}({\bm{x}}^{(j)}(t))$ that are $\bm{u}$}-coordinates of the trajectories $\{{\bm{x}}^{(j)}(t):t\geq 0\}$, $j=1,\dots,J$, and ${\bm{U}}=\text{proj}({\bm{X}})$
@@
    Find
    $(j,s)=\operatorname*{argmin}_{l\in\{1,\dots,J\},t\geq 0}\|{\bm{U}}-{\bm{u}}^{(l)}(t)\|$ (16)
@@
    Define $G:\mathbb{R}^{n}\to\{\gamma^{(j)}:j=1,\dots,J\}$, where
    $G({\bm{X}})={\bm{x}}^{(j)}(s).$ (17)
We chose to use the Euclidean distance $\|\cdot\|$ in (16) for its simplicity and speed of computation, but the method can be applied using other distance metrics. Figure 5 illustrate steps (5-7) above.
Figure 5: Phase correction. The phase correction $G({\bm{X}}_{i})$ for a point ${\bm{X}}_{i}$ on the stochastic trajectory is computed in three steps. First, the RRE solution and point ${\bm{X}}_{i}$ is projected into the $\bm{u}$}-coordinates of the non-linear Centre manifold. Then, the point, ${\bm{u}}(s_{i})$ lying on the projected RRE trajectory $\{{\bm{u}}(t):t\geq 0\}$ that minimises the distance to ${\bm{U}}_{i}$ is found. Finally, the corresponding point, ${\bm{x}}(s_{i})$, with the same time-point $s_{i}$ with ${\bm{u}}(s_{i})$ and that lies on the RRE solution $\{{\bm{x}}(t):t\geq 0\}$ is found. The point ${\bm{x}}(s_{i})$ and the perturbation $\bm{\kappa}_{i}$ (instead of ${\bm{x}}_{i}$ and $\bm{\xi}_{i}$) are used in the next LNA transition. The magnitude of $\bm{\kappa}_{i}$ is smaller or equal than the magnitude of $\bm{\xi}_{i}$ on the $\bm{u}$}-coordinates.
@@ 3 Hopf bifurcation systems
Condition (H.I) is precisely the requirement for the macroscopic reaction rate equation to exhibit an equilibrium. Condition (H.II) ensures the equilibrium is non-hyperbolic and therefore prescribes the existence of a parameter dependent, two-dimensional, attracting manifold near $({\bm{x}}_{\alpha_{p}},\alpha_{p})$ that is l.t.e. to a system of the form as in (15). Furthermore, the complex pair of eigenvalues of the Jacobian matrix implies that, near $({\bm{x}}_{\alpha_{p}},\alpha_{p})$, the system is l.t.e. to the Generic Hopf bifurcation system (see Figure 4).

### v4

@@ 3.2 The pcLNA algorithm
The Algorithm 1 uses the standard LNA transition in steps 1-2 interrupted by phase-correcting maps in steps 3-4. The map $G$ is used to find the phase for ${\bm{X}}(t_{i})$. After phase correction, the phase of the system is moved from $s_{i-1}+\Delta t$ to $s_{i}$, the deterministic state moved from ${\bm{x}}^{(j_{i-1})}(s_{i-1}+\Delta t)$ to ${\bm{x}}^{(j_{i})}(s_{i})$ and the fluctuations $\bm{\xi}(s_{i-1}+\Delta t)$ are replaced with the remaining fluctuations after phase correction, $\bm{\kappa}(s_{i})$. Provided $\Delta t$ is not too large, phase drifts occurring in the next LNA transition will be small. Indeed, $\{\bm{\kappa}(s_{i}):i\in\mathbb{N}_{0}\}$ should have uniformly bounded variance and remain sufficiently small for the LNA predictions to be valid over the short time intervals $s_{i},s_{i}+\Delta t),\ i\in\mathbb{N}_{0}$. Important for computational efficiency is that the same solutions of ODEs (6), (12) and (13) are used in all steps. The fact that these ODEs only need to be solved once before any simulations, rapidly increases the computation speed. The steps are illustrated in Figure 3.
Figure 3: General pcLNA algorithm over steps i1,i,i+1i-1,i,i+1. For simplicity we illustrate the case where j1=j=j+1j-1=j=j+1, so the same RRE solution is used over these steps. pcLNA proceeds with standard LNA steps interrupted by phase corrections with ${\bm{x}}^{(j_{i})}(t_{i})$ replaced by $G({\bm{X}}(t_{i}))={\bm{x}}^{(j_{i})}(s_{i})$, and the perturbations $\bm{\xi}(t_{i})$ replaced by $\bm{\kappa}(s_{i})$.
All that is preventing one from readily applying the steps in the pcLNA algorithm is an explicit phase-defining metric $d$ and a set of reference trajectories $\Gamma$ used in the phase-correcting map G{G} as in (15), (16). A natural question arises: does there exist a single choice of metric $d$ and reference set $\Gamma$ that can be used for all reaction networks to produce accurate simulations over long times? As we will explain, the answer is no. Phase depends on the system’s underlying dynamics, and in particular on which components generate persistent non-linear behavior. Consequently, no universal metric can correctly capture phase across qualitatively different systems.
In the next section, we show how tools from dynamical systems theory provide a way to determine a phase-defining metric $d$ and the set of reference trajectories $\Gamma$ tailored to the system’s qualitative dynamics.
3.3 Choosing the phase-defining metric $d$ and set of reference trajectories $\Gamma$
We use three connected concepts from the dynamical systems theory to guide our approach in choosing metric $d$ and set $\Gamma$ used in the phase-correcting map $G$. The first concept is topological equivalence, which allows us to group systems exhibiting qualitatively identical dynamics. The second concept is decomposition, enabling us to break down complex (potentially high-dimensional) systems into a (possibly low-dimensional) non-linear component with persistent dynamics and components with transient dynamics. The third concept involves exemplar systems within each group that are well understood and can serve as references. We highlight the aspects of these concepts that are used in our approach in Sections 3.3.1, 3.3.2 and 3.3.3 and explain how we practically apply them in the construction of suitable metric $d$ and set $\Gamma$ in Section 3.3.4. We will then use this approach in Sections 4 and 5 to define $d$ and $\Gamma$ for systems presenting a Hopf bifurcation and a different $d$ and $\Gamma$ for systems exhibiting bistability through a fold bifurcation.
3.3.1 Topological equivalence
Two continuous dynamical systems of the form (6) are topologically equivalent if their dynamics are qualitatively identical. Formally, this holds if there exists a homeomorphism $\bm{h}$} which maps paths of one system onto paths of the either while preserving the direction of time. Local topological equivalence (l.t.e.) occurs when this holds in neighborhoods of the systems’ domains (see Definitions 2.14-5 in [37], or SI Section S3).
A practical approach to identify l.t.e. is to examine the eigenvalues of the Jacobian in (7) near an equilibrium point $({\bm{x}}_{\alpha_{p}},\alpha_{p})$, with ${\bm{F}}({\bm{x}}_{\alpha_{p}},\alpha_{p})={\bm{0}}$. Let $_{0}, n+n_{+}$, and $n_{-}$ be the number of eigenvalues of $\bm{J}({\bm{x}}_{\alpha_{p}},\alpha_{p})$ that have zero, positive, and negative real part, respectively. An equilibrium is hyperbolic if $n_{0}=0$; otherwise, it is non-hyperbolic. For hyperbolic equilibria, l.t.e. is ensured if $n_{+}$ and $n_{-}$ match between systems (see Theorem 2.2 in [37]); the non-hyperbolic case is more subtle and discussed below.
3.3.2 Decomposition
Here we derive a novel decomposition of a stochastic process following the LNA to coordinates with different type of dynamics using results from the centre manifold theory (see [37, 59, 62, 51]), and Ito’s differential calculus (see [38]). For a detailed description of the results used in this section and the calculations performed see SI Section S3.
Suppose the system (6) exhibits an equilibrium $({\bm{x}}_{\alpha},\alpha)$ which is non-hyperbolic ($n_{0}>0$) when $\alpha=\alpha_{p}$. In this setting, the system’s non-linear dynamics can be reduced to occur on an $\alpha$-dependent $n_{0}$-dimensional centre manifold tangent to the eigenspace corresponding to the eigenvalues of $\bm{J}({\bm{x}}_{\alpha_{p}},\alpha_{p})$ with zero real part (see Theorem 5.1 and Lemma 5.1 in [37]).
In fact, the system is l.t.e. to
$\left\{\begin{aligned} \dot{{\bm{u}}}&=\bm{f}({\bm{u}},\alpha),\\ \dot{{\bm{v}}}&={\bm{A}}_{\text{s}}({\bm{u}},\alpha){\bm{v}},\\ \dot{\bm{w}}&={\bm{A}}_{\text{u}}({\bm{u}},\alpha)\bm{w},\end{aligned}\right$. (18)
for paths and parameter values near to $({\bm{x}}_{\alpha_{p}},\alpha_{p})$ where for each $t\geq 0$, ${\bm{u}}(t)\in\mathbb{R}^{n_{0}}$, ${\bm{v}}(t)\in\mathbb{R}^{n_{-}}$, $\bm{w}(t)\in\mathbb{R}^{n_{+}}$, $f:\mathbb{R}^{n_{0}}\times\mathbb{R}\to\mathbb{R}^{n_{0}}$ is smooth, ${\bm{A}}_{\text{s}}({\bm{u}}(t),\alpha)$ is an $n_{-}\times n_{-}$ (stable) matrix with all of its eigenvalues having negative real part, and ${\bm{A}}_{\text{u}}({\bm{u}}(t),\alpha)$ is an $n_{+}\times n_{+}$ (unstable) matrix with all of its eigenvalues having positive real part (See Section 4 in [62] or Section 2 in [51]). Moreover, if the respective local homeomorphism denoted by ${\bm{h}}_{\alpha}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ satisfies a set of generic nonresonance conditions of the eigenvalues of $\bm{J}({\bm{x}}_{\alpha_{p}},\alpha_{p})$ that hold almost surely (see SI, S3), ${\bm{h}}_{\alpha}$ is at least twice differentiable in a neighborhood of ${\bm{x}}_{\alpha_{p}}$. In this setting, Itô’s lemma (see [38, Theorem 5.10]) can be applied locally to characterise the image of the LNA under ${\bm{h}}_{\alpha}$:
$\begin{pmatrix}{\bm{U}}(t)\\ {\bm{V}}(t)\\ {\bm{W}}(t)\end{pmatrix}:={\bm{h}}_{\alpha}({\bm{x}}(t)+\Omega^{-1/2}\bm{\xi}(t))=\begin{pmatrix}{\bm{u}}(t)\\ {\bm{v}}(t)\\ \bm{w}(t)\end{pmatrix}+\Omega^{-1/2}\begin{pmatrix}\bm{\xi}_{{\bm{u}}}(t)\\ \bm{\xi}_{{\bm{v}}}(t)\\ \bm{\xi}_{\bm{w}}(t)\end{pmatrix}+O(\Omega^{-1}),$ (19)
for times $t\geq 0$ where ${\bm{x}}(t)+\Omega^{-1/2}\bm{\xi}(t)$ remains close to ${\bm{x}}_{\alpha_{p}}$. Here $({\bm{u}}(t),{\bm{v}}(t),\bm{w}(t))^{\top}:={\bm{h}}_{\alpha}({\bm{x}}(t))$ evolves according to (18) and the perturbation $\tilde{\bm{\xi}}(t)=(\bm{\xi}_{{\bm{u}}}(t),\bm{\xi}_{{\bm{v}}}(t),\bm{\xi}_{\bm{w}}(t))^{\top}$ evolves according to the system of SDEs
$d\tilde{\bm{\xi}}=\tilde{\bm{J}}_{t}\tilde{\bm{\xi}}dt+\tilde{\bm{E}}_{t}dB_{t}$ (20)
where
$\tilde{J}_{t}=\begin{bmatrix}\nabla_{\bm{u}}\bm{f}({\bm{u}},\alpha)&0&0\\ \nabla_{\bm{u}}({\bm{A}}_{\text{s}}({\bm{u}},\alpha){\bm{v}})&{\bm{A}}_{\text{s}}({\bm{u}},\alpha)&0\\ \nabla_{\bm{u}}({\bm{A}}_{\text{u}}({\bm{u}},\alpha)\bm{w})&0&{\bm{A}}_{\text{u}}({\bm{u}},\alpha)\end{bmatrix},\quad\tilde{\bm{E}}_{t}=\nabla_{\bm{x}}{\bm{h}}_{\alpha}({\bm{x}}(t)){\bm{E}}_{t}$ (21)
Importantly under this local homeomorphism, the LNA, omitting $O(\Omega^{-1})$ terms, can be decomposed into 3 processes:
@@ 3.3.2 Decomposition
    ${\bm{U}}(t)\approx{\bm{u}}(t)+\Omega^{-1/2}\bm{\xi}_{\bm{u}}(t)$ the sum of a non-linear deterministic process, and a stochastic process with drift $\nabla_{\bm{u}}\bm{f}({\bm{u}},\alpha)\bm{\xi}_{u}$;
@@
    ${\bm{V}}(t)\approx{\bm{v}}(t)+\Omega^{-1/2}\bm{\xi}_{\bm{v}}(t)$ the sum of an exponentially decaying deterministic process and a linear stochastic process with stable drift matrix,
@@
    ${\bm{W}}(t)\approx\bm{w}(t)+\Omega^{-1/2}\bm{\xi}_{\bm{w}}(t)$ the sum of an exponentially growing deterministic process and a linear stochastic process with unstable drift matrix.
That is, at the $O(\Omega^{-1/2})$ scale, ${\bm{V}}(t)$ quickly drifts toward zero and ${\bm{W}}(t)$ quickly drifts outside the neighborhood, leaving all the effective stochastic behavior of the LNA to be described by ${\bm{U}}(t)$. This implies that any phase drift responsible for the long-term failure of the LNA arises from the dynamics of ${\bm{U}}(t)$ relative to ${\bm{u}}(t)$; the other processes, ${\bm{V}}(t)$ and ${\bm{W}}(t)$, play no long-term role and do not contribute to persistent phase drift.
Throughout the rest of the paper, we refer to the first $n_{0}$ coordinates of the image of ${\bm{h}}_{\alpha}$, where ${\bm{U}}(t)$ and ${\bm{u}}(t)$ evolve, as the centre coordinates; the next $n_{-}$ coordinates, where ${\bm{V}}(t)$ and ${\bm{v}}(t)$ evolve, as the stable coordinates; and last $n_{+}$ coordinates, where ${\bm{W}}(t)$ and $\bm{w}(t)$ evolve, as the unstable coordinates.
While this mapping is generally non-linear, its action on the centre coordinates admits a first-order approximation near the equilibrium (See SI Section S3). Specifically, for any point $\bm{X}$} sufficiently close to ${\bm{x}}_{\alpha_{p}}$, its location $\bm{U}$} in the centre coordinates of ${\bm{h}}_{\alpha}({\bm{X}})$ satisfies
${\bm{U}}=\bm{P}_{c}({\bm{X}}-{\bm{x}}_{\alpha_{p}})+O(\lVert{\bm{X}}-{\bm{x}}_{\alpha_{p}}\rVert^{2}),$ (22)
where $\bm{P}_{c}$ denotes the projection onto the eigenspace of $\bm{J}({\bm{x}}_{\alpha_{p}},\alpha_{p})$ corresponding to eigenvalues of zero real part. It is important to note this is a first-order approximation of these coordinates and valid only sufficiently close to the equilibrium. Consequently, given a realization of the LNA, ${\bm{X}}(t)={\bm{x}}(t)+\Omega^{-1/2}\bm{\xi}(t)$, the locations of ${\bm{U}}(t)$ are approximately
${\bm{U}}(t)\approx\bm{P}_{c}({\bm{X}}(t)-{\bm{x}}_{\alpha_{p}}),$ (23)
where $\bm{P}_{c}={\bm{R}}_{c}^{-1}$ and ${\bm{R}}_{c}$ is the $n\times n_{0}$ matrix whose columns form an eigenbasis corresponding to the eigenvalues of $\bm{J}({\bm{x}}_{\alpha_{p}},\alpha_{p})$ with zero real part. Similarly,
${\bm{u}}(t)\approx\bm{P}_{c}({\bm{x}}(t)-{\bm{x}}_{\alpha_{p}}).$ (24)
3.3.3 Normal forms
Normal forms are simple exemplar systems representing classes of l.t.e. systems. Low-dimensional centre manifolds have well-studied normal forms, including the fold and Hopf bifurcations (see [37], Figure 4). We use normal forms in two ways. First, to specify the conditions under which a given system is l.t.e. to the normal form, typically related to the eigen-decomposition of the Jacobian. Secondly, we use their simpler form to identify the key qualitative features (e.g., oscillations, multi-stability) shared by systems in their class, guiding the choice of an appropriate metric $d$ and set $\Gamma$ for the phase-correcting map $G$.
3.3.4 Procedure for constructing the phase-correcting map
The theory in Sections 3.3.1-3.3.3 lead us to the following principles for defining the phase-defining metric $d$ and set of reference trajectories $\Gamma$:
  1. A.
    Local topological equivalence (l.t.e.) identifies classes of systems whose trajectories are qualitatively identical near equilibria allowing the same construction of the phase-defining metric $d$ and reference set $\Gamma$ to be applied across all systems in the class.
  2. B.
    Decomposition isolates the coordinates in which non-linear, persistent dynamics — and hence phase drifts — occur. Phase and phase correction is defined exclusively in these centre coordinates, while the transient dynamics of the stable and unstable coordinates are disregarded when constructing the metric $d$.
  3. C.
    Normal forms provide canonical low-dimensional representatives of each l.t.e. class and make explicit the geometric structure of trajectories in the centre coordinates. This structure determines both the appropriate phase-defining metric and the choice of representative reference trajectories $\Gamma$.
Using these principles we define a procedure for constructing the phase-correcting map $G$ for any system as in (6).
  1. 1.
    Check if there exists a parameter value $\alpha_{p}$ and a state ${\bm{x}}_{\alpha_{p}}$ in a neighborhood of interest such that ${\bm{F}}({\bm{x}}_{\alpha_{p}},\alpha_{p})={\bm{0}}$, and the Jacobian $\bm{J}({\bm{x}}_{\alpha_{p}},\alpha_{p})$ has $n_{0}>0$ eigenvalues with zero real part. If so, proceed to the next step; otherwise, consider using an alternative method (including the standard LNA).
  2. 2.
    Identify the local topological equivalence (l.t.e.) class of (6) by checking whether it satisfies the conditions of a known normal form, such as those associated with Hopf bifurcations or bi-stable systems (see Sections 4 and 5).
  3. 3.
    Based on the l.t.e. class dynamics and, where applicable, the corresponding normal form, choose a set of reference trajectories $\Gamma=\bigcup_{\ell=1}^{L}\{{\bm{x}}^{(\ell)}(t):t\geq 0\}$, containing at least one deterministic solution of each distinct type of qualitative behaviour. For example, if the system admits three stable equilibria, select three trajectories converging to each equilibrium.
@@ 3.3.4 Procedure for constructing the phase-correcting map
    Define a projection map to find the approximate location of a point near ${\bm{x}}_{\alpha_{p}}$ in the centre coordinates:
    $\text{proj}:\mathbb{R}^{n}\to\mathbb{R}^{n_{0}},\quad\text{proj}({\bm{x}})={\bm{R}}_{c}^{-1}({\bm{x}}-{\bm{x}}_{\alpha_{p}}),$ (25)
    where ${\bm{R}}_{c}$ is an $n\times n_{0}$ matrix whose columns form an eigenbasis corresponding to the eigenvalues of $\bm{J}({\bm{x}}_{\alpha_{p}},\alpha_{p})$ that have zero real part.
@@
    Compute the projections of all reference trajectories
    $\{{\bm{u}}^{(\ell)}(t)=\text{proj}({\bm{x}}^{(\ell)}(t)):t\geq 0\},\quad\ell\in\mathcal{L},\,\,\mathcal{L}=\{1,\dots,L\}$ (26)
    and, for any stochastic state ${\bm{X}}\in\mathbb{R}^{n}$, define
    ${\bm{U}}:=\text{proj}({\bm{X}}).$ (27)
@@
    Define the distance metric $d:\mathbb{R}^{n}\times\mathbb{R}^{n}\to[0,\infty)$, by
    $d({\bm{y}},{\bm{x}}):=\|\text{proj}({\bm{y}})-\text{proj}({\bm{x}})\|.$ (28)
@@
    Using (25)-(28), the phase-correcting map $G:\mathbb{R}^{n}\to\Gamma$ defined in (15) is given by
    $G({\bm{X}})={\bm{x}}^{(j)}(s),\text{ where }(j,s)=\operatorname*{argmin}_{\ell\in\mathcal{L},\,t\geq 0}d({\bm{X}},{\bm{x}}^{(\ell)}(t))=\operatorname*{argmin}_{\ell\in\mathcal{L},\,t\geq 0}\|{\bm{U}}-{\bm{u}}^{(\ell)}(t)\|.$ (29)
In summary, this procedure defines the phase-correcting function $G$, which maps any stochastic state $\bm{X}$} to the point on the reference trajectories in $\Gamma$ that is closest to $\bm{X}$} in the approximate centre coordinates. Figure 5 illustrates steps (5-8) above.
Although the true centre coordinates are given by the $\alpha$-dependent map ${\bm{h}}_{\alpha}$, the linear projection (25) provides a first-order approximation that is independent of $\alpha$ and can therefore be used consistently across parameter values.
For the numerical investigations in Section 6, we use the Euclidean norm (28), though other metrics may be employed.
In Section 4 and 5 we apply this procedure for two concrete types of dynamics: Oscillations (sustained or damped) and bistability which occur when the RRE in (6) contains a Hopf bifurcation and generic fold bifurcation respectively. We show the resulting metrics $d$ and sets $\Gamma$ can be used to accurately simulate such dynamics in Section 6.
Figure 5: Phase correction. The phase-correcting map $G({\bm{X}}_{i})$ for a point ${\bm{X}}_{i}$ on the stochastic trajectory is computed in three steps. First, the RRE solution and point ${\bm{X}}_{i}$ is projected into the centre coordinates of the non-linear Centre manifold. Then, the point, ${\bm{u}}(s_{i})$ lying on the projected RRE trajectory $\{{\bm{u}}(t):t\geq 0\}$ that minimises the distance to ${\bm{U}}_{i}$ is found. Finally, the corresponding point, ${\bm{x}}(s_{i})$, with the same time-point $s_{i}$ with ${\bm{u}}(s_{i})$ and that lies on the RRE solution $\{{\bm{x}}(t):t\geq 0\}$ is found. The point ${\bm{x}}(s_{i})$ and the perturbation $\bm{\kappa}_{i}$ (instead of ${\bm{x}}_{i}$ and $\bm{\xi}_{i}$) are used in the next LNA transition. The magnitude of $\bm{\kappa}_{i}$ is smaller than or equal to the magnitude of $\bm{\xi}_{i}$ on the centre coordinates.
@@ 4 Hopf bifurcation systems
Condition (H.I) is precisely the requirement for the macroscopic reaction rate equation to exhibit an equilibrium. Condition (H.II) ensures the equilibrium is non-hyperbolic and therefore prescribes the existence of a parameter dependent, two-dimensional, attracting manifold near $({\bm{x}}_{\alpha_{p}},\alpha_{p})$ for which (6) is l.t.e. to a system
[…]
where for each $t\geq 0$, ${\bm{u}}(t)\in\mathbb{R}^{2},\ {\bm{v}}(t)\in\mathbb{R}^{n-2},\ \bf:\mathbb{R}^{2}\times\mathbb{R}\to\mathbb{R}^{2}$ is smooth and ${\bm{A}}({\bm{u}}(t),\alpha)$ is an $(n-2)\times(n-2)$ matrix with all of its eigenvalues having negative real part. See Figure 4 for an illustration of the paths of the normal form of this system.
Since all the eigenvalues other than the complex pair have negative real values, then the calculations in Section 3.3.2 reveal the image of the LNA, ${\bm{X}}(t)={\bm{x}}(t)+\Omega^{-1/2}\bm{\xi}(t)$, under the local homeomorphism reveals two stochastic processes: ${\bm{U}}(t)\approx{\bm{u}}(t)+\Omega^{-1/2}\bm{\xi}_{\bm{u}}(t)$, evolving in the centre coordinates and exhibiting the non-linear dynamics; and ${\bm{V}}(t)\approx{\bm{v}}(t)+\Omega^{-1/2}\bm{\xi}_{\bm{v}}(t)$, evolving in the stable coordinates and decaying at an exponential rate. Their stochastic perturbations, $\bm{\xi}_{\bm{u}}(t)$ and $\bm{\xi}_{\bm{v}}(t)$, evolve according to (20).